Here's the question:
If the radius of certain circle is tripled, with 8.2 cm then added, the result is the circumference of the circle. Find the radius of the circle.
Ok I am assuming part of this is 2(pi)r= r^3 + 8.2 but then I get stuck!! Please help!! Thank you!

Answer by stanbon(75887)

If the radius of certain circle is tripled, with 8.2 cm then added, the result is the circumference of the circle. Find the radius of the circle.
---------------------------------
EQUATION:
3r+8.2 = 2(pi)r
r(3-2pi) = -8.2
r = -8.2/(3-2pi)
r = 2.49757...cm
=====================
Cheers,
Stan H.